$SPAD/src/input egg.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

draw(surface(5*sin(u)*cos(v),4*sin(u)*sin(v),3*cos(u)),u=0..%pi,v=0..2*%pi)

———-

References^{[1] nothing}

4